OpenStudy (anonymous):

Find the inverse of x/sqrtx^2+7

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x ^{2}+7}}\]

OpenStudy (anonymous):

seems unlikely that this is a one to one function, so it will not have an inverse

OpenStudy (anonymous):

I know, it looks weird but apparently there is an inverse

OpenStudy (anonymous):

apparently the inverse is \[\frac{ \sqrt{7x} }{ \sqrt{1-x^{2}} }\]

OpenStudy (anonymous):

well i assure you it does not on the other hand you can write \[x=\frac{y}{\sqrt{y^2+7}}\] and solve for \(y\)

OpenStudy (anonymous):

we can do it if you like

OpenStudy (anonymous):

that's what I did

OpenStudy (anonymous):

still couldn't find it

OpenStudy (anonymous):

nevermind, I got it lol

OpenStudy (anonymous):

i skipped a step in my head so i couldn't find out what i did wrong on the paper

OpenStudy (anonymous):

\[x=\frac{y}{\sqrt{y^2+7}}\] \[x\sqrt{y^2+7}=y\] square both sides \[x^2(y^2+7)=y^2\] distribute \[x^2y^2+7x^2=y^2\] put all terms with \(y\) on one side, all else on the other \[x^2y^2-y^2=-7x^2\] factor \[y^2(x^2-1)=-7x^2\] divide \[y^2=\frac{-7x^2}{x^2-1}\]

OpenStudy (anonymous):

or if you prefer \[y^2=\frac{7x^2}{1-x^2}\] but now how are you going to solve that for \(y\)??

OpenStudy (anonymous):

yeah i see it now, thanks for the help

OpenStudy (anonymous):

square root both sides

OpenStudy (anonymous):

you get \[y=\pm\sqrt{\frac{7x^2}{1-x^2}}\] which is certainly not a function

OpenStudy (anonymous):

don't forget the \(\pm\)

OpenStudy (anonymous):

well you could actually get \[\frac{ \sqrt{7x} }{ \sqrt{1-x ^{2}} }\] which is a function